1. An experiment was run where varying amounts of HClO and KClO solutions were mixed and the pH of the resulting solution was measured. A plot of the data is given below.
Y = 1*X + 7.53
^^^ a graph with pH on the y-axis and log(KClO/HCLO) on the x-axis

A. What is the conjugate base in this mixture?
B. What is the acid in this mixture?
C. What is the Ka value for this acid?
D. If you have a 0.100 mol/L solution of this acid, what is the pH of the solution?
E. If you have a 0.0100 mol/L solution of this conjugate base, what is the pH of the solution?
Answer #1

(A) Conjugate base is ClO-.

(B) Acid is HClO

(C) When [KClO] = [HClO] , the corresponding pH is the pKa of the acid. In that condition, log( KClO/ HClO) = 0

pKa = 0 + 7.53 = 7.53

So, Ka = 10-7.53 = 2.95 * 10-8

(D) When, [HClO] = 0.100 M

So, [H+] = ( Ka * [HClO]) 1/2 = ( 0.100 * 2.95 * 10-8 )1/2 = 5.43 * 10-5 M

So, pH = - log [H+] = - log (5.43 * 10-5 ) = 4.26